### v2

@@ 3. Data-Centric Subgraph Isomorphism Algorithm
In this section, we introduce a data-science-inspired method for solving the subgraph isomorphism problem. Instead of exhaustively searching for full pattern matches in the data graph via node-by-node traversal, we adopt a modular strategy based on motif decomposition. We call this approach $\Delta$-Motif due to the decomposition of the pattern graphs into smaller components. In Section 4, we show an analogy of $\Delta$-Motif to the popular $\Delta$-Stepping algorithm used for SSSP (Single-Source Shortest Path) problems (meyer2003delta, 40). Although the two algorithms address entirely different problems, both methods illustrate a similar trade-off accepting potentially additional work in exchange for higher performance. Yet, a detailed discussion of this analogy will only become apparent after a thorough discussion and performance analysis of our new $\Delta$-Motif algorithm.
@@ 3.1. Preliminaries
Figure 1. (a) Heavy-hex and (b) 2D square grid lattice structures used as data graphs in our benchmark tests. These lattices represent common qubit connectivity topologies in quantum hardware and and serve as the input data graphs for quantum layout selection. The corresponding set of representative motifs is shown in (c).
@@
We denote the set of all embeddings of a motif $M_{i}$ in the data graph by $Res(M_{i})$. Each element in $Res(M_{i})$ is a tuple of vertex indices that together form a subgraph isomorphic to $M_{i}$. These embeddings are stored in tabular form, where each row represents a single match and each column corresponds to a vertex position in the motif template. This tabular format allows subsequent processing, such as joining embeddings and enforcing constraints, to be performed using scalable dataframe operations.
@@ 3.2. The $\Delta$-Motif Approach
To make this process concrete, we now present a step-by-step example of finding all subgraph isomorphisms. As a first step, we select a set of motifs to be used to decompose $G_{p}$. The number of unique motifs can explode with increasing graph size, and thus creating a list of all possible motifs is infeasible and thus we limit ourselves to a small subset of these. The motif set $\mathcal{M}$ can either be predefined or adaptively selected based on the characteristics of the data graph. The efficiency of the overall decomposition process depends heavily on this choice. Larger motifs reduce the number of decompositions needed but are more expensive to compute and store, while smaller motifs offer greater flexibility and lower overhead. We discuss detailed motif selection strategies in later sections, and for now, focus on demonstrating the workflow.
@@
Our decomposition records the overlapping vertices between slices, which later serve as the constraints in the join phase. In this example, the overlaps occur at $v_{2}$ and $v_{5}$, corresponding to the constraints $S_{0}[0]=S_{1}[3]$ and $S_{0}[3]=S_{1}[0]$. We now proceed to join the motif embeddings corresponding to the slices. For clarity, we denote the two dataframes as $Res(M_{4})_{\text{L}}$ and $Res(M_{4})_{\text{R}}$, representing the left and right sides of the join operation. Based on the overlapping nodes identified during decomposition, the join constraints are defined as $Res(M_{4})_{\text{L}}[0]==Res(M_{4})_{\text{R}}[3]$ and $Res(M_{4})_{\text{L}}[3]==Res(M_{4})_{\text{R}}[0]$, where the bracket notation refers to entire columns of vertex assignments at the respective motif embeddings. These constraints are used as join keys in the operation, which is efficiently implemented in standard dataframe libraries and relational database systems. The join step eliminates invalid combinations and retains only those pairs of embeddings that satisfy the required boundary conditions. Fig. 2(e) illustrates the join operation and its output to clarify the overall process.
Although the join operation enforces the specified boundary constraints between motif embeddings, it may still produce combinations with unintended vertex overlaps. These arise when two embeddings share additional vertices beyond the expected overlaps, violating the one-to-one vertex mapping required for valid subgraph isomorphism. To address this, we introduce a filter step that removes such invalid combinations. Fig. 2(f) illustrates an example of this issue. In general, the number of duplicated vertices in each joined row should exactly match the number of join constraints. Any row containing additional duplications is deemed invalid and is discarded. This is similar to the pruning phase of a tree-search.
@@ 3.3. Pattern Graph Decomposition Using Motifs
In principle many motifs could be used for decomposing the graph, we typically restrict ourselves to a small, carefully-chosen family that balances coverage and computational cost. The choice of motifs has a significant impact on the performance of the algorithm. Larger motifs do not necessarily guarantee faster execution, as their embeddings $Res(M)$ often result in more expensive join operations. Including more motifs in the set also increases the upfront preparation time. In later section 6.3, we discuss how to select an appropriate motif set based on the characteristics of both the data and pattern graphs, and support these design choices with empirical results.
@@
While this may seem counterintuitive as we replace one subgraph isomorphism problem with another, we highlight the following: 1) instead of performing subgraph isomorphism on the data graph, we apply it to the typically much smaller pattern graph, 2) the motifs we seek are typically orders of magnitude smaller than both the data and pattern graph, and 3) in our motif decomposition, we only require finding a single match of the motif in the pattern, allows us to terminate the search early rather than exhaustively enumerating all possible matches. Under these conditions, the tree-search based algorithm is actually more suitable and efficient for this task.
Once a slice $S_{i}$ is identified, the algorithm inspects all vertices in $S_{i}$ to determine which have outgoing edges that remain uncovered in the current pattern graph. These vertices are designated as boundary nodes, since they will serve as connection points for subsequent slices. For example, in Section 3.2, vertices $v_{2}$ and $v_{5}$ from $S_{0}$ play this role. The non-boundary nodes from $S_{i}$ are removed from the pattern graph to form a reduced graph for the next iteration, while the boundary nodes are retained to ensure overlap with subsequent matches. This procedure is repeated in the following iterations. Each new slice is matched using the reduced pattern graph, and its overlapping vertices with the union of previously selected slices are recorded. These overlapping vertices define structural dependencies across slices and are stored as join constraints. These constraints will later be used in the join step to reassemble the full pattern structure.
@@ 3.5. Performance Increase from Larger Motifs
The use of larger motifs for decomposing the pattern graph has several benefits: 1) fewer motifs are needed for the decomposition, hence fewer join-and-filter , which can be substantially fewer than the branching levels explored in a VF2 tree search, 2) larger motifs can increase parallelism in the data pre-processing phase, and 3) larger motifs enable early-stage pruning of invalid solutions, which also takes place during the pre-processing phase.
In our performance analysis, we show that using larger motifs will outperform the $M_{2}$-only case. Consider the case where a pattern graph contains an $M_{9}$ path. While we can find the matches of $M_{9}$ from successive joins of many $M_{2}$ motifs, requiring multiple iterations of merging and filtering. In contrast, the same path can be obtained in a single step by joining two larger embeddings, such as $Res(M_{5})\Join Res(M_{5})$.
@@ 4. Analog To Tree Search And Prune
Now that we have presented our new join-based algorithm for subgraph isomorphism, we can draw a parallel to the traditional tree-search and prune-based approaches. While not apparent at first glance, these algorithms share quite a few algorithmic features but also trade off a few.
@@
Recall that most tree-search and prune-based solutions rely on recursion. While some parallelism is possible, it is mostly attainable by creating an ordered set of vertices that represent the first few levels of the tree search (e.g., one to three) and parallelizing across those ordered elements. Extending parallelism beyond these initial levels can present several challenges: workload imbalance, larger memory footprint for storing starting points (similar to generating $Res(M_{3})$), and most critically, synchronization overhead as parallel threads must coordinate access to shared data structures. For these reasons, the tree search and prune algorithms face inherent limitations in the degree of parallelization that they can achieve.
In contrast, most phases of $\Delta$-Motif are inherently parallel because each row in our dataframe represents an independent candidate. The algorithm applies the same operations uniformly across all rows, enabling massive parallelization where thousands of candidate mappings can be processed simultaneously. This parallelism comes at a cost: our new algorithm has an increased memory footprint. The increased memory requirements stem from two main factors: 1) holding intermediate data generated during join operations before key constraints are applied to prune invalid matches and 2) storing complete tables of intermediate solutions after each join-and-filter iteration.
@@ 4.2. $M_{2}$ and Its Analogy to Parallel Tree-Search
Consider the case where the pattern graph is decomposed only with the $M_{2}$ motif. In this setup, the number of steps in $\Delta$-Motif matches the branching levels explored in a VF2 tree search. The difference is that VF2 moves forward one branch at a time in sequence, while $\Delta$-Motif processes all branches at a given level in parallel. For example, the first $M_{2}$ expansion enumerates all edges, the next expands them into all possible 3-hop paths, then into 4-hop paths, and so on. This mirrors the depth expansion of VF2 but instead of walking the tree branch by branch, $\Delta$-Motif resolves each level through joins and filters that run simultaneously. This approach may require more intermediate storage, but the payoff is the ability to handle entire layers of the search space at once, effectively trading storage for scalability and providing a parallel counterpart to VF2’s sequential tree-search-and-prune. The results presented in the following sections show that this trade-off yields significant performance gains.
$M_{2}$-only case share the similar inefficiencies as tree-search based approaches like VF2. They must still explore all shorter paths of length 2, 3, and 4, pruning non-relevant branches along the way. This recursion incurs significant overhead, and the same path may be traversed multiple times from different starting points. It also implies that we might pay the penalty of traversing the bad path multiple times. In contrast, when using larger motif choices, $\Delta$-Motif does not have that problem. Lastly, the irregular branching patterns in tree search causes CPUs to struggle due to branch mispredictions. Such mispredictions have been shown to be costly in both sorting (green2014merging, 41) and graph algorithms (greenbranchavoiding, 42), with the irregular traversals of VF2 proving to be equally challenging for hardware to predict.
@@ 4.3. $\Delta$-Motif and $\Delta$-Stepping
With all the above now in place, we want to draw one last analogy to the $\Delta$-Stepping algorithm used for solving the SSSP problem. And while, there is no similarity between SSSP and subgraph isomorphism, there is an interesting connection between the two approaches. In the case of $\Delta$-Stepping, the parameter $\Delta$ acts as a tuning knob that allows trading off parallelization with computational efficiency (close to theoretical bounds). Larger $\Delta$ values allow vertices to be traversed earlier, at the cost of potentially revisiting them multiple times. The key emphasis is that by using the $\Delta$-Stepping technique we can expose parallelization at the expense of of additional memory. Similarly, our $\Delta$-Motif approach also exposes parallelization at multiple granularities: across multiple paths all within the same motif, across the entire join-and-filter operations, and across the preparation of all the motif embeddings in pre-processing computation. Together, these factors enable $\Delta$-Motif to achieve strong performance on both CPUs and GPUs.
@@ 5.1. Software and Hardware Configurations
One of the goals of this work is to demonstrate that the subgraph isomorphism could be implemented using data-science tools, allowing for straightforward deployments across CPU and GPU while maintaining high performance. To achieve this, we intentionally avoid implementing specialized kernels, but relied on using open-source and commodity frameworks. Moreover, our implementation is designed to be flexible enough to run efficiently across hardware platforms.
@@
Our CPU implementation leverages both Pandas (Pandas, 43) and NumPy (NumPy, 44), representing motif embeddings and candidate isomorphisms as Pandas DataFrames. We rely on Pandas’ efficient merge operations and NumPy’s vectorized filter capabilities for core processing. While these libraries do not use multi-threading by default, they achieve performance gains through SIMD (Single Instruction Multiple Data) execution (numpy_simd, 45).
@@
$\Delta$-Motif on the GPU is implemented using cuPy (Cupy2017, 16) and cuDF 25.6.0 (cudf-24.02, 17), which serve as drop-in replacements for NumPy and Pandas, respectively. Due to strong interoperability across the NVIDIA RAPIDS ecosystem and adherence to Apache Arrow standards, our algorithm can transition between CPU and GPU execution with only minimal changes to Python import statements. While subgraph isomorphism is a graph problem, our implementation does not rely on cuGraph or NetworkX (NetworkX, 46). Instead, it directly operates on tabular data representations using cuDF and cuPy. We note that RAPIDS provides a rich suite of additional GPU-accelerated libraries, ncluding cuML, cuGraph, and cuSpatial, that can be leveraged in broader workflows.
@@
We use the VF2 algorithm as our baseline, representing the broader class of tree-search-and-prune methods for subgraph isomorphism. Introduced over two decades ago (Cordella2004, 10), VF2 remains a widely used standard in both research and production, with mature implementations available in NetworkX (NetworkX, 46), the Boost Graph Library (boost_graph_library, 47), and rustworkx(Rustworkx2022, 14). For our experiments, we selected rustworkx, a high-performance Rust implementation that maintains a NetworkX-compatible API. Unlike NetworkX’s pure Python version, which is easy to use but limited in performance(raamurri2024comparative, 48), rustworkx delivers up to 100× speedup and is used in production systems such as IBM’s Qiskit quantum computing framework (qiskit2024, 49). This ensures that our comparisons reflect true algorithmic differences rather than implementation bottlenecks.

### v3

@@ 3. Data-Centric Subgraph Isomorphism Algorithm
In this section, we introduce a data-science-inspired method for solving the subgraph isomorphism problem. Instead of exhaustively searching for full pattern matches in the data graph via node-by-node traversal, we adopt a modular strategy based on motif decomposition. We call this approach $\Delta$-Motif, reflecting the decomposition of pattern graphs into smaller components. In Section 4, we draw an analogy between $\Delta$-Motif and the popular $\Delta$-Stepping algorithm for single-source shortest path (SSSP) problems (undefaai, ), highlighting a similar trade-off between additional work and improved parallelism. The details of this analogy become clearer after introducing the algorithm and its performance characteristics.
@@ 3.1. Preliminaries
Figure 1. (a) Heavy-hex and (b) 2D square grid lattice structures used as data graphs in our benchmark tests. These lattices represent common qubit connectivity topologies in quantum hardware and serve as the input data graphs for quantum layout selection. The corresponding set of representative motifs is shown in (c).
@@
We denote the set of all embeddings of a motif $M_{i}$ in the data graph by $Res(M_{i})$. Each element in $Res(M_{i})$ is a tuple of vertex indices that together form a subgraph isomorphic to $M_{i}$. These embeddings are stored in tabular form, where each row corresponds to a single match and each column to a vertex position in the motif template. This tabular format allows subsequent processing, such as joining embeddings and enforcing constraints, to be performed using scalable dataframe operations.
@@ 3.2. The $\Delta$-Motif Approach
To make this process concrete, we now present a step-by-step example of finding all subgraph isomorphisms. As a first step, we select a set of motifs to be used to decompose $G_{p}$. The number of unique motifs can explode with increasing graph size, and thus creating a list of all possible motifs is infeasible. Thus, we limit ourselves to a small subset of these. The motif set $\mathcal{M}$ can be predefined or adaptively selected based on the characteristics of the data graph. The efficiency of the overall decomposition process depends heavily on this choice. Larger motifs reduce the number of necessary decompositions, but are more expensive to compute and store. Whereas smaller motifs offer greater flexibility and lower overhead. We discuss detailed motif selection strategies in later sections, and for now, we focus on demonstrating the workflow.
@@
Our decomposition records the overlapping vertices between slices, which later serve as constraints in the join phase. In this example, the overlaps occur at $v_{2}$ and $v_{5}$, corresponding to the constraints $S_{0}[0]=S_{1}[3]$ and $S_{0}[3]=S_{1}[0]$. We now proceed to join the motif embeddings corresponding to the slices. For clarity, we denote the two dataframes as $Res(M_{4})_{\text{L}}$ and $Res(M_{4})_{\text{R}}$, representing the left and right sides of the join operation. Based on the overlapping nodes identified during decomposition, the join constraints are defined as $Res(M_{4})_{\text{L}}[0]==Res(M_{4})_{\text{R}}[3]$ and $Res(M_{4})_{\text{L}}[3]==Res(M_{4})_{\text{R}}[0]$, where the bracket notation refers to entire columns of vertex assignments at the respective motif embeddings. These constraints are used as join keys in the operation, which is efficiently implemented in standard dataframe libraries and relational database systems. The join step eliminates invalid combinations and retains only those pairs of embeddings that satisfy the required boundary conditions. Fig. 2(e) illustrates the join operation and its output to clarify the overall process.
@@
Although the join operation enforces the specified boundary constraints between motif embeddings, it may still produce combinations with unintended vertex overlaps. These arise when two embeddings share additional vertices beyond the expected overlaps, violating the one-to-one vertex mapping required for valid subgraph isomorphism. To address this, we introduce a filter step that removes such invalid combinations. Fig. 2(f) illustrates an example of this issue. In general, the number of duplicated vertices in each joined row should exactly match the number of join constraints. Any row containing additional duplicates is deemed invalid and discarded. This is similar to the pruning phase of a tree-search.
@@ 3.3. Pattern Graph Decomposition Using Motifs
In principle, many motifs could be used for decomposing the graph; we typically restrict ourselves to a small, carefully-chosen family that balances coverage and computational cost. The choice of motifs has a significant impact on the performance of the algorithm. Larger motifs do not necessarily guarantee faster execution, as their embeddings $Res(M)$ often result in more expensive join operations. Including more motifs in the set also increases the upfront preparation time. In section 6.3, we discuss how to select an appropriate motif set based on the characteristics of both the data and pattern graphs, and support these design choices with empirical results.
@@
While this may seem counterintuitive as we replace one subgraph isomorphism problem with another, we highlight the following: 1) instead of performing subgraph isomorphism on the data graph, we apply it to the typically much smaller pattern graph, 2) the motifs we seek are typically orders of magnitude smaller than both the data and pattern graph, and 3) in our motif decomposition, we only require finding a single match of the motif in the pattern, allows us to terminate the search early rather than exhaustively enumerating all possible matches. Under these conditions, the tree-search based algorithm is actually more suitable and efficient for this task. Lastly, the motif decomposition searches for a decomposition rather than the ideal decomposition.
Once a slice $S_{i}$ is identified, the algorithm inspects all vertices in $S_{i}$ to determine which have outgoing edges that remain uncovered in the current pattern graph. These vertices are designated as boundary nodes. Boundary nodes serve as connection points for subsequent slices. For example, in Section 3.2, vertices $v_{2}$ and $v_{5}$ from $S_{0}$ play this role. The non-boundary nodes from $S_{i}$ are removed from the pattern graph to form a reduced graph for the next iteration, while the boundary nodes are retained to ensure overlap with subsequent matches. This procedure is repeated in the following iterations. Each new slice is matched using the reduced pattern graph, and its overlapping vertices with the union of previously selected slices are recorded. These overlapping vertices define structural dependencies across slices and are stored as join constraints. These constraints will later be used in the join step to reassemble the full pattern structure.
@@ 3.5. Performance Increase from Larger Motifs
The use of larger motifs for decomposing the pattern graph has several benefits: 1) fewer motifs are needed for the decomposition, hence fewer join-and-filter , which can be substantially fewer than the branching levels explored in a VF2 tree search, 2) larger motifs can increase parallelism in the data pre-processing phase, and 3) larger motifs enable early-stage pruning of invalid solutions, which also takes place during the pre-processing phase. In our performance analysis, we show that using larger motifs will outperform the $M_{2}$-only case.
@@ 4. Analog To Tree Search And Prune
Now that we have presented our new join-based algorithm for subgraph isomorphism, we can draw a parallel to the traditional tree-search and prune-based approaches. While not apparent at first glance, these algorithms share quite a few algorithmic features but also trade off a few. The literature is full of algorithms that alternate between BFS and DFS for conducting the pattern search.
@@
Recall that most tree-search and prune-based solutions rely on recursion. While some parallelism is possible, it is mostly attainable by creating an ordered set of vertices that represent the first few levels of the tree search (e.g., one to three) and parallelizing across those ordered elements – this one example in which a BFS is conducted at the beginning to generate initial starting conditions and is then followed by DFS operations that focus on finding the remaining pattern. Extending parallelism beyond these initial levels can present several challenges: workload imbalance, larger memory footprint for storing starting points (similar to generating $Res(M_{3})$), and most critically, synchronization overhead as parallel threads must coordinate access to shared data structures. For these reasons, the tree search and prune algorithms face inherent limitations in the degree of parallelization that they can achieve. And this also explains why some algorithms are limited to finding patterns with fewer than 10 vertices and edges.
In contrast, most phases of $\Delta$-Motif are inherently parallel because each row in our dataframe represents an independent candidate. The algorithm applies the same operations uniformly across all rows, enabling massive parallelization where thousands of candidate mappings can be processed simultaneously – similar to the BFS operation. This parallelism comes at a cost: our new algorithm has an increased memory footprint. The increased memory requirements stem from two main factors: 1) holding intermediate data generated during join operations before key constraints are applied to prune invalid matches and 2) storing complete tables of intermediate solutions after each join-and-filter iteration.
@@ 4.2. $M_{2}$ and Its Analogy to Parallel Tree-Search
Consider the case where the pattern graph is decomposed only with the $M_{2}$ motif. In this setup, the number of steps in $\Delta$-Motif matches the branching levels explored in a VF2 tree search. The difference is that VF2 moves forward one branch at a time in sequence, while $\Delta$-Motif processes all branches at a given level in parallel (similar to a BFS traversal). For example, the first $M_{2}$ expansion enumerates all edges, the next expands them into all possible 3-hop paths, then into 4-hop paths, and so on. This mirrors the depth expansion of VF2 but instead of walking the tree branch by branch, $\Delta$-Motif resolves each level through joins and filters that run simultaneously. This approach may require more intermediate storage, but the payoff is the ability to handle entire layers of the search space at once, effectively trading storage for scalability and providing a parallel counterpart to VF2’s sequential tree-search-and-prune. The results presented in the following sections show that this trade-off yields significant performance gains.
The $M_{2}$-only case exhibits inefficiencies similar to tree-search approaches such as VF2, as it must still explore shorter paths of length 2, 3, and 4 while pruning non-relevant branches along the way. This recursion incurs significant overhead, and the same path may be traversed multiple times from different starting points. It also implies that we might pay the penalty of traversing the bad path multiple times. In contrast, using larger motif choices allows $\Delta$-Motif to operate on larger structural units, avoiding repeated traversal; invalid candidates are discarded during filtering rather than through recursive backtracking. Additionally, the irregular branching patterns inherent in tree-search algorithms lead to frequent branch mispredictions, which have been shown to be costly in both sorting (undefaaj, ) and graph algorithms (undefaak, ). The irregular traversal patterns of VF2 exhibit similar challenges for modern processors.
@@ 4.3. $\Delta$-Motif and $\Delta$-Stepping
We now return to the analogy with the $\Delta$-Stepping algorithm for solving the SSSP problem. Although SSSP and subgraph isomorphism address different problems, both approaches expose additional parallelism by allowing extra work during execution. In $\Delta$-Stepping, the parameter $\Delta$ trades computational efficiency for parallelism by allowing vertices to be processed earlier at the cost of potentially revisiting them. Similarly, our $\Delta$-Motif exposes parallelism at multiple granularities: across multiple paths all within the same motif, across the entire join-and-filter operations, and across the preparation of all the motif embeddings in pre-processing computation. Together, these design choices enable $\Delta$-Motif to achieve strong performance on both CPUs and GPUs.
@@ 5.1. Software and Hardware Configurations
One of the goals of this work is to demonstrate that the subgraph isomorphism could be implemented using open-source and well-defined data-science frameworks and standard APIs, enabling execution on both CPU and GPU platforms without specialized kernels while maintaining high performance. Moreover, our implementation is designed to be flexible enough to run efficiently across hardware platforms.
@@
Our CPU implementation leverages both Pandas (undefaal, ) and NumPy (undefaam, ), representing motif embeddings and candidate isomorphisms as Pandas DataFrames. We rely on Pandas’ efficient merge operations and NumPy’s vectorized filter capabilities for core processing. While these libraries do not use multi-threading by default, they achieve performance gains through Single Instruction Multiple Data (SIMD) execution (undefaan, ).
@@
$\Delta$-Motif on the GPU is implemented using cuPy (undefr, ) and cuDF 25.6.0 (undefs, ), which serve as drop-in replacements for NumPy and Pandas, respectively. Due to strong interoperability across the NVIDIA RAPIDS ecosystem and adherence to Apache Arrow standards, our algorithm can transition between CPU and GPU execution with only minimal changes to Python import statements. While subgraph isomorphism is a graph problem, our implementation does not rely on cuGraph or NetworkX (undefaao, ). Instead, it directly operates on tabular data representations using cuDF and cuPy. We note that RAPIDS provides a rich suite of additional GPU-accelerated libraries, including cuML, cuGraph, and cuSpatial, that can be leveraged in broader workflows.
@@
We use the VF2 algorithm as our baseline, representing the broader class of tree-search-and-prune methods for subgraph isomorphism. Introduced over two decades ago (undefi, ), VF2 remains a widely used standard in both research and production, with mature implementations available in NetworkX (undefaao, ), the Boost Graph Library (undefaap, ), and rustworkx(undefp, ). For our experiments, we selected rustworkx, a high-performance Rust implementation that maintains a NetworkX-compatible API. Unlike NetworkX’s pure Python version, which is easy to use but limited in performance(undefaaq, ), rustworkx delivers up to 100× speedup and is used in production systems such as IBM’s Qiskit quantum computing framework (undefaar, ). This ensures that our comparisons reflect true algorithmic differences rather than implementation bottlenecks.
GSI
We also include GPU-friendly Subgraph Isomorphism (GSI) (undefae, ; undefaas, ) as a representative GPU-based baseline. GSI is a widely used GPU-oriented subgraph matching algorithm designed to improve parallel efficiency through GPU-centric execution strategies. In our evaluation, we report GSI results on benchmarks involving real-world network datasets with small pattern graphs, such as triangle enumeration, where the reference implementation is applicable. For experiments involving larger pattern graphs, GSI performance results were not included as the solution did not complete within reasonable execution time under our experimental settings or gave runtime errors due to out-of-memory.